Luis is tiling the rectangular floor of a room measuring 8 feet 6 inches by 12 feet. How many 6-inch-by-6-inch tiles will Luis need to tile the floor without overlapping? Also, I would like to know how you answer the problem not just the answer. Thank you!
Answer by jim_thompson5910(35256)

8 ft 6 in = 8ft + 6in = 96 in + 6 in = 102 in


Now divide this by 6 to get: 102/6 = 17

So he needs 17 tiles along the width.

12 ft = 144 inches

144/6 = 24

So he needs 24 tiles along the length.

Therefore, he needs 17*24 = 408 tiles total